OpenStudy (anonymous):

If x=5a, what is the average (arithmetic mean) of x and a in terms of a? A. 2a B. 3a C. 4a D. 5a E. 6a

OpenStudy (anonymous):

add and divide by 2

OpenStudy (anonymous):

\[\frac{x+a}{2}=\frac{5a+a}{2}\] is a start

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

can you finish from there?

OpenStudy (anonymous):

3a ?